OpenStudy (marigirl):

Can you please explain.. integrate 2 cos^2(x)

OpenStudy (anonymous):

probs a good idea to use trig identities

OpenStudy (anonymous):

\[\cos(2x)=2\cos ^{2}(x)-1\]

OpenStudy (marigirl):

so then i have: 2 cos^2(x)=cos(2x)+1

OpenStudy (marigirl):

when i integrate that I get sin(2x)/2 + x....... the ans says sin(x)cos(x)+x

OpenStudy (anonymous):

use trig identities for sin(2x)

OpenStudy (anonymous):

sin(2x) = 2sin(x)cos(x)

OpenStudy (marigirl):

thanks!

OpenStudy (anonymous):

: )

OpenStudy (anonymous):

http://www.purplemath.com/modules/idents.htm

OpenStudy (anonymous):

probs a good idea to have a cheat sheet by your side when you do these problems

OpenStudy (anonymous):

cheat sheet of trig identities...

OpenStudy (marigirl):

even the cheat sheet is complicated :(

OpenStudy (anonymous):

not if you know what you are looking for! if you find it hard to integrate something and its not your basic cosx sin(x) etc...then most likely its a trig identity you need to use to replace it

OpenStudy (marigirl):

or i just need to do more problems! haha

OpenStudy (anonymous):

yep, repitition and u will be a gun

OpenStudy (marigirl):

thanks for that website link!

OpenStudy (anonymous):

np : )